# DyBBT: Dynamic Balance via Bandit inspired Targeting for Dialog Policy with Cognitive Dual-Systems

### v1

@@ B.5.2 Knowledge Distillation Buffer Management
B.5.3 Visitation Count of the Cognitive State Space
@@
This choice of dimensions is motivated by cognitive and dialog theory, which highlights stage, uncertainty, and structural relationships as key factors influencing decision making. By quantifying these environmental affordances into a structured cognitive state space $\mathcal{C}$, we create a formal bridge between Gibson’s ecological perception theory and practical dialog policy optimization. While not exhaustive, this representation aims to capture the most salient features for guiding exploration. Its empirical necessity and sufficiency are validated through ablation studies in Section 4.3. We define $\mathcal{C}$ as the cognitive state space, assumed to be a compact subset of $\mathbb{R}^{3}$ equipped with the Euclidean metric $d(\mathbf{c},\mathbf{c}^{\prime})$.
B.5.4 Calculation of Empirical Cumulative Regret
To empirically validate the theoretical intuition of sublinear regret growth under our simplifying assumptions, we compute the empirical cumulative regret $R_{\text{emp}}(T)$ during training, as shown in Figure 7. The regret is defined as:
@@ B.5.4 Calculation of Empirical Cumulative Regret
$R_{\text{emp}}(T)$ is ploted against $T$ on a log-log scale to clearly visualize the sublinear growth trend. The theoretical upper bound $\widetilde{\mathcal{O}}(\sqrt{T})$ is plotted alongside for comparison. The constant factor in the theoretical bound is fit to the empirical curve in the early training phase to align the curves for illustrative purposes.
@@ Appendix C Human Evaluation Details
This appendix provides comprehensive details of the human evaluation study described in Section 4.4. The study was designed to qualitatively assess the core contribution of the DyBBT framework: the intelligent, adaptive decision making of its meta-controller, beyond what is captured by automated metrics.
@@ C.1 Annotation Protocol and Interface
    Action Appropriateness: “How appropriate is the system’s chosen action given the current dialog state?" Rated on a 5 points Likert scale:
@@
    Switching Judgment: “In this specific situation, would it be justified to invoke a powerful, but computationally expensive, reasoning module to choose the action?" Answered with Yes or No. This question was only shown for states where the evaluated model did not invoke System 2, to directly test if the meta-controller’s decision not to invoke aligned with human judgment.
@@ C.3 Human Evaluation Results
Model Variant Action Appropriateness $\uparrow$ Switching Agreement $\uparrow$
DyBBT-8B 4.31 $\pm$ 0.12 88.7%
w/o Meta-Controller (Random) 3.72 $\pm$ 0.19 52.3%
w/ S1-only 3.95 $\pm$ 0.15
w/o Exploration Condition (EC) 4.08 $\pm$ 0.14 75.4%
w/o Confidence Condition (CC) 3.89 $\pm$ 0.16 81.2%
@@ C.4 Qualitative Analysis of Meta-Controller Decisions
Belief State:
restaurant {
semi {
food: "Chinese" # (USER_CONFIRMED)
pricerange: "cheap" # (USER_CONFIRMED)
area: "" # (USER_MENTIONED but NOT_CONFIRMED)
name: "" # (NOT_MENTIONED - High Uncertainty)
book { people: "", day: "", time: "" }
taxi { ... } # (Not relevant in this turn)
Listing 1: Belief state exemplifying high user uncertainty.
Cognitive State: $d_{t}=0.3$ (early-stage), $u_{t}=0.8$ (high uncertainty), $\rho_{t}=0.6$. Meta-Controller Decision: System 1’s confidence was low ($p_{t}^{S1}=0.6<\kappa$), triggering System 2 via the confidence condition. System 2 performed a multi path reasoning and produced a confirm_all action sequence to disambiguate the user’s intent: confirm(restaurant, area) and confirm(restaurant, name). Annotators overwhelmingly rated this intervention as appropriate (Avg: 4.8/5) and agreed (90%) that invoking System 2 was justified. This case demonstrates the critical role of the confidence condition as a robustness safeguard against System 1’s inherent limitations in partially observable contexts.
Case 2: Low Agreement Example (Random Switching). A random switching baseline ($10\%$ chance per turn) invoked System 2 in a state where the optimal action was obvious, leading to computational waste without performance gain:
Belief State:
restaurant {
semi {
food: "Chinese" # (CONFIRMED)
pricerange: "cheap" # (CONFIRMED)
area: "east" # (CONFIRMED)
name: "Golden Dragon" # (CONFIRMED)
book {
people: "4", day: "today", time: "19:00" # (BOOKED)
taxi {
semi {
departure: "train station", # (CONFIRMED)
destination: "Golden Dragon", # (CONFIRMED)
leaveAt: "19:30" # (CONFIRMED)
Listing 2: Belief state where the task is complete.
Cognitive State: $d_{t}=0.9$ (late stage), $u_{t}=0.1$ (low uncertainty), $\rho_{t}=0.2$. Scenario: All user constraints are satisfied, and the booking is complete. The only appropriate action is to terminate the dialog with goodbye. The random controller invoked System 2, which also output goodbye. Annotators rated the action itself as appropriate (Avg: 4.2/5) but unanimously (100%) judged the invocation of System 2 as not justified, deeming it an inefficient use of resources. This highlights a key failure mode of static or non-adaptive switching heuristics and underscores the necessity of our cognitive state aware meta-controller.
In summary, these cases provide concrete evidence that DyBBT’s switching mechanism dynamically allocits computational resources in a manner that is both effective and efficient, closely mirroring human expert judgment.
Appendix D Further Experimental Analysis
D.1 Experimental Results on MultiWOZ
Table 4 presents DyBBT’s performance on the MultiWOZ multi domain dialog dataset, including key metrics (Inform, Success, Book, Turns). Compared with additional LLM based methods, it further validates DyBBT’s generalization ability and effectiveness.
Table 4: Evaluation results on MultiWOZ dataset. DyBBT-8B/GPT-4.0 denotes Qwen3-8B for System 1 and GPT-4.0 for System 2. DQN, LLM_DP and EIERL are reported in EIERLZhao et al. (2025), other results were reported from original papers, “—" indicates unreported results.
@@ D.1 Experimental Results on MultiWOZ
D.2 Ablation Study Settings and Results
@@
Table 5: Ablation study on MultiWOZ.
[…]
DyBBT-8B (full) 91.2 84.1 86.9 14.6
w/o Meta-Controller 82.5 71.8 77.3 17.5
w/o System 2 85.7 76.3 80.1 16.8
w/ Learned Cognitive State 90.5 83.2 86.3 14.8
w/o Knowledge Distillation 89.8 82.4 85.7 15.1
w/o Cognitive State (raw $s_{t}$) 84.2 75.1 79.6 17.1
w/o Exploration Condition (EC) 90.1 82.9 86.1 14.9
w/o Confidence Condition (CC) 87.6 79.5 83.2 16.2
w/o dialog Progress ($d_{t}$) 88.9 80.7 84.5 15.7
w/o User Uncertainty ($u_{t}$) 89.6 81.9 85.3 15.3
w/o Slot Dependency ($\rho_{t}$) 90.3 82.5 85.9 15.0
D.3 Supplementary Analysis Figures and Tables
@@
This subsection provides all supplementary figures and tables supporting the main text analysis in Section 5, which offer intuitive data support for the discussions:
@@
  • Figure 5: Heatmap of visitation frequency in the cognitive state space $\mathcal{C}$, illustrating the structured exploration strategy of the meta-controller across dialog phases.
@@
  • Figure 5: Analysis of meta-controller decisions, showing the rate of System 2 invocation across dialog progress and the proportion of triggers from each condition.
@@
  • Figure 7: Demonstrates the improvement of System 1 through knowledge distillation and the corresponding reduction in System 2 invocation over training.
@@
  • Figure 7: Compares the empirical cumulative regret of DyBBT against the theoretical upper bound derived under simplifying assumptions.
@@
D.4 Hyperparameter Sensitivity Analysis
A key concern is the sensitivity of DyBBT’s performance to the meta-controller’s hyperparameters: the exploration threshold $\tau$, the confidence threshold $\kappa$, and the number of bins used to discretize the cognitive state space $\mathcal{C}$. We conducted a comprehensive grid search over $\tau\in\{0.5,1.0,1.5,2.0\}$, $\kappa\in\{0.5,0.6,0.7,0.8,0.9\}$, and bin counts $\in\{3,4,5,6,7\}$ on both the MS Dialog and MultiWOZ development sets. Performance is measured by the success rate (%), and the results are visualized in Figure 8.

### v2

@@ B.5.3 Knowledge Distillation Buffer Management
B.5.4 Visitation Count of the Cognitive State Space
@@
This choice of dimensions is motivated by cognitive and dialog theory, which highlights stage, uncertainty, and structural relationships as key factors influencing decision making. By quantifying these environmental affordances into a structured cognitive state space $\mathcal{C}$, we create a formal bridge between Gibson’s ecological perception theory and practical dialog policy optimization. While not exhaustive, this representation aims to capture the most salient features for guiding exploration. Its empirical necessity and sufficiency are validated through ablation studies in Sec. 4.3. We define $\mathcal{C}$ as the cognitive state space, assumed to be a compact subset of $\mathbb{R}^{3}$ equipped with the Euclidean metric $d(\mathbf{c},\mathbf{c}^{\prime})$.
B.5.5 Calculation of Empirical Cumulative Regret
To empirically validate the theoretical intuition of sublinear regret growth under our simplifying assumptions, we compute the empirical cumulative regret $R_{\text{emp}}(T)$ during training, as shown in Fig. 7. The regret is defined as:
@@ B.5.5 Calculation of Empirical Cumulative Regret
$R_{\text{emp}}(T)$ is plotted against $T$ on a log-log scale to clearly visualize the sublinear growth trend. The theoretical upper bound $\widetilde{\mathcal{O}}(\sqrt{T})$ is plotted alongside for comparison. The constant factor in the theoretical bound is fit to the empirical curve in the early training phase to align the curves for illustrative purposes.
@@ Appendix C Human Evaluation Details
This appendix provides comprehensive details of the human evaluation study described in Sec. 4.4. The study was designed to qualitatively assess the core contribution of the DyBBT framework: the intelligent, adaptive decision making of its meta-controller, beyond what is captured by automated metrics.
@@ C.1 Annotation Protocol and Interface
    Action Appropriateness: “How appropriate is the system’s chosen action given the current dialog state?” Rated on a 5 points Likert scale:
@@
    Switching Judgment: “In this specific situation, would it be justified to invoke a powerful, but computationally expensive, reasoning module to choose the action?” Answered with Yes or No. This question was only shown for states where the evaluated model did not invoke System 2, to directly test if the meta-controller’s decision not to invoke aligned with human judgment.
@@ C.3 Human Evaluation Results
Model Variant Action Appropriateness $\uparrow$ Switching Agreement $\uparrow$ DyBBT-8B 4.31 $\pm$ 0.12 88.7% w/o Meta-Controller (Random) 3.72 $\pm$ 0.19 52.3% w/ S1-only 3.95 $\pm$ 0.15 w/o Exploration Condition (EC) 4.08 $\pm$ 0.14 75.4% w/o Confidence Condition (CC) 3.89 $\pm$ 0.16 81.2%
@@ C.4 Qualitative Analysis of Meta-Controller Decisions
Cognitive State: $d_{t}=0.3$ (early-stage), $u_{t}=0.8$ (high uncertainty), $\rho_{t}=0.6$. Meta-Controller Decision: System 1’s confidence was low ($p_{t}^{S1}=0.6<\kappa$), triggering System 2 via the confidence condition. System 2 performed a multi path reasoning and produced a confirm_all action sequence to disambiguate the user’s intent: confirm(restaurant, area) and confirm(restaurant, name). Annotators overwhelmingly rated this intervention as appropriate (Avg: 4.8/5) and agreed (90%) that invoking System 2 was justified. This case demonstrates the critical role of the confidence condition as a robustness safeguard against System 1’s inherent limitations in partially observable contexts. Case 2: Low Agreement Example (Random Switching). A random switching baseline ($10\%$ chance per turn) invoked System 2 in a state where the optimal action was obvious, leading to computational waste without performance gain: Cognitive State: $d_{t}=0.9$ (late stage), $u_{t}=0.1$ (low uncertainty), $\rho_{t}=0.2$. Scenario: All user constraints are satisfied, and the booking is complete. The only appropriate action is to terminate the dialog with goodbye. The random controller invoked System 2, which also output goodbye. Annotators rated the action itself as appropriate (Avg: 4.2/5) but unanimously (100%) judged the invocation of System 2 as not justified, deeming it an inefficient use of resources. This highlights a key failure mode of static or non-adaptive switching heuristics and underscores the necessity of our cognitive state aware meta-controller. In summary, these cases provide concrete evidence that DyBBT’s switching mechanism dynamically allocates computational resources in a manner that is both effective and efficient, closely mirroring human expert judgment.
Appendix D Real World User Experiments
Table 5: Real World User Experiment Results. Success Rate measures the percentage of successfully completed dialogs. Average Turns counts the number of dialog turns per task. User Satisfaction is rated on a 1-5 Likert scale.
@@ Appendix D Real World User Experiments
While all previous experiments relied on simulated users, real world user interactions are inherently more complex and unpredictable. This raises a key concern regarding generalization: user behavior in practice may not neatly align with the quantifiable dimensions of our cognitive state space $\mathcal{C}$, potentially limiting DyBBT’s applicability. To investigate this and verify the robustness of our assumptions, we conducted experiments with real human users.
D.1 Experimental Settings and Analysis
We recruited 30 volunteers with natural language interaction experience, each completing 10 sets of multi-domain dialogs. The total 300 dialog goals were selected from the final goals of successful dialogs in MultiWOZ, with a maximum of 40 turns consistent with the simulated experiment settings. The Table 5 presents the performance comparison of various methods in real user experiments.
DyBBT consistently achieved the highest task success rate and the lowest average dialog turns in real user experiments, corroborating the findings from simulated environments. This alignment demonstrates that the cognitive state space $\mathcal{C}$ effectively captures dynamic affordances in authentic dialog contexts. Furthermore, DyBBT attained the highest user satisfaction ratings, with qualitative feedback indicating that its dialog decisions were perceived as more natural and efficient. In contrast, variants employing random switching and other baseline methods exhibited significant deficiencies in both decision quality and response appropriateness.
Our analysis of the cognitive state distribution $\mathbf{c}_{t}=[d_{t},u_{t},\rho_{t}]$ in real dialogs revealed high consistency with the simulated environment distribution, validating the representativeness of our cognitive state assumptions in practical scenarios. The majority of failure cases were attributable to user intent shifts, irrelevant user responses, or non-standard linguistic input, challenges inherent to real world dialog systems that highlight areas for future improvement.
D.2 Case Studies
To provide qualitative insights into DyBBT’s performance with real users, we present three representative dialog cases from our experiments.
D.2.1 Case 1: Mid-Dialog User Intent Shift
The user initially requested a “Chinese food” restaurant reservation but suddenly asked “Are there any attraction nearby?” midway through the dialog.
Cognitive State:
  • $d_{t}=0.4$ (mid-stage)
  • $u_{t}=0.6$ (original goal incomplete; new intent introduces uncertainty)
  • $\rho_{t}=0.3$ (low dependency; cross domain intent)
DyBBT’s Decision:
  • System 1’s confidence $p_{t}^{S1}=0.55<\kappa$, triggering System 2.
  • System 2 generated multi-path strategies: 1.Confirm whether to continue the original task; 2.Provide attraction information; 3.Confirm to end the original task and ask attraction information.
Outcome: System 2 selected the “confirmation first” path, outputting ‘confirm(restaurant, food, “Chinese”)’ → ‘inform(attraction, type, “cinema”)’. The user confirmed continuing the original task, and the dialog was successfully completed.
Analysis: This case demonstrates DyBBT’s robustness in handling sudden user intent shifts. Although $u_{t}$ and $\rho_{t}$ in the cognitive state did not directly reflect the severity of the intent conflict, System 1’s low confidence accurately captured the uncertainty of the current strategy, thereby triggering System 2’s in-depth reasoning. Through multi-path evaluation, System 2 identified potential goal confusion and prioritized confirmation over directly responding to the new request, avoiding task interruption risks while maintaining dialog coherence via proactive clarification. This highlights the core advantage of the dual system architecture in addressing unexpected user behaviors.
D.2.2 Case 2: Vague User Information
When requesting a taxi, the user said “I want to go to that famous square” without specifying the exact name.
Cognitive State:
  • $d_{t}=0.2$ (early stage)
  • $u_{t}=0.9$ (high uncertainty)
  • $\rho_{t}=0.7$ (strong dependency between destination and departure location)
DyBBT’s Decision:
  • The cognitive state region had a low visitation count ($n_{t}(\mathbf{c}_{t})=8<\tau\sqrt{\log T}$), triggering System 2.
  • System 2 generated strategies: 1.Request the exact name; 2.List possible options; 3.Confirm the area.
Outcome: System 2 selected the “option list” path, outputting ‘request(taxi, destination)’ → ‘inform(attraction, name, “Central Square”)’. The user made a selection, and the task proceeded.
Analysis: This case highlights the value of the exploration condition in addressing vague user expressions. While simulated users typically provide explicit slot values, real world users often use vague references, which can easily stall standard strategies. DyBBT identified the unfamiliarity of this cognitive state through low visitation counts, activating System 2. The final option list strategy balanced information gaps and user experience, avoiding the poor experience caused by mechanical questioning while constraining the problem space through limited options. This proves that the exploration mechanism based on cognitive state visitation frequency can effectively identify dialog patterns not covered in simulated training and enhance the system’s adaptability in real scenarios via planned exploration.
D.2.3 Case 3: Non-Typical User Behavior
After completing a hotel reservation, the user suddenly repeatedly asked “Does the price include breakfast?”.
Cognitive State:
  • $d_{t}=0.9$ (late stage)
  • $u_{t}=0.1$ (low uncertainty; all slots confirmed)
  • $\rho_{t}=0.2$ (low dependency)
DyBBT’s Decision:
  • The cognitive state was judged as “well explored” with high System 1 confidence ($p_{t}^{S1}=0.88$), so System 2 was not triggered.
  • System 1 directly responded with ‘inform(hotel, breakfast, “no”)’.
Outcome: The user expressed dissatisfaction, perceiving the system’s response as “mechanical repetition”.
Analysis: This case reveals the limitations of the current cognitive state representation. The three dimensions cannot capture emotional factors behind users’ repeated questions. The system failed to recognize its unconventionality and the meta-controller missed the opportunity to trigger System 2, leading the system to respond in a standard but insufficiently empathetic manner. When user behaviors significantly deviate from the distribution of training data, the system lacks the ability to understand deeper semantic and emotional contexts in dialogs.
Appendix E Further Experimental Analysis
E.1 Experimental Results on MultiWOZ
Table 6 presents DyBBT’s performance on the MultiWOZ multi-domain dialog dataset, including key metrics (Inform, Success, Book, Turns). Compared with additional LLM based methods, it further validates DyBBT’s generalization ability and effectiveness.
[…]
Table 6: Evaluation results on MultiWOZ dataset. DyBBT-8B/GPT-4.0 denotes Qwen3-8B for System 1 and GPT-4.0 for System 2. DQN, LLM_DP and EIERL are reported in EIERLZhao et al. (2025), other results were reported from original papers, “—” indicates unreported results.
E.2 Ablation Study Settings and Results
@@
E.3 Confidence Condition Error Analysis
To further clarify the crucial role of the Confidence Condition (CC) in the DyBBT framework, we conducted an in depth analysis of the types and proportions of errors prevented by this mechanism. The CC primarily serves as a safety net to prevent System 1 from making “catastrophic errors” in states with “high cognitive uncertainty”, whereas the absence of the Exploration Condition (EC) mainly leads to reduced “exploration efficiency” rather than direct task failures.
E.3.1 Types and Proportions of Errors Prevented by the CC
We analyzed a Sample of 200 CC interventions dialog logs of "DyBBT w/o EC" and "DyBBT w/o CC". Table 7 summarizes the distribution of error types among these cases.
Error Type Description Proportion Impact Level 1. Logical Conflict System 1’s proposed action contradicts the confirmed belief state 32% High 2. Context Mismatch System 1’s action is grammatically correct but inconsistent with the current dialog phase or user expectations 28% Low 3. Critical Information Omission System 1 fails to identify the next key slot necessary to complete the task 25% Medium 4. Domain/Slot Confusion System 1 confuses slots or selects the wrong domain in cross domain scenarios 15% High
Table 7: Types and proportions of errors prevented by the Confidence Condition
Types 1 and 4 account for 47% of errors, which are relatively severe and would almost certainly lead to dialog failure if not corrected by the CC. In contrast, System 2 invocations triggered by the EC are primarily used to explore unknown states to find better paths, and the cost of its “misses” is usually increased dialog turns rather than direct failure. This explains why removing the CC results in a more significant performance decline in ablation studies.
E.3.2 Case Analysis of CC Interventions
The following four cases demonstrate how the CC prevents serious errors in practice.
Case 1: CC Prevents a “Logical Conflict” Error
  • Background: After the user booked a restaurant, they requested a taxi.
  • Belief State:
  • System 1’s Incorrect Action: request(taxi, destination) with confidence $p_{t}^{S1}=0.62$ (low confidence, triggering CC)
  • CC’s Intervention: System 2 identified the logical flaw: the destination should be the restaurant, and the departure location should be queried instead. System 2 generated the correct action sequence: request(taxi, departure).
  • Analysis: System 1 made a fundamental logical error. Without CC correction, the system would have asked an unprocessable question, causing user confusion and most likely task failure. The CC acts as a commonsense reasoning validator here.
Case 2: CC Prevents a “Context Mismatch” Error
  • Background: Late in the dialog, all user constraints were satisfied, and the hotel was successfully booked.
  • Belief State:
  • System 1’s Suboptimal Action: request(hotel, parking) with confidence $p_{t}^{S1}=0.58$ (low confidence, triggering CC)
  • CC’s Intervention: System 2 determined that continuing to request information after task completion was redundant and confusing. The optimal action was to politely end the dialog. System 2 output: inform(booking, complete) $\rightarrow$ goodbye().
  • Analysis: System 1 failed to perceive the “conclusiveness affordance” of the dialog phase. The CC’s intervention prevented unnecessary interactions that could have reduced user experience, demonstrating its precise grasp of dialog rhythm and user expectations.
Case 3: CC Prevents a “Critical Information Omission” Error
  • Background: The user is booking a restaurant, having provided the number of people and date but not the time.
  • Belief State:
  • System 1’s Incorrect Action: inform(restaurant, name, "Pasta Place") with confidence $p_{t}^{S1}=0.55$ (low confidence, triggering CC)
  • CC’s Intervention: System 2 reasoned that the necessary booking time must be obtained before providing the restaurant name; otherwise, the booking process cannot be completed. System 2 generated the action sequence: request(restaurant, time).
  • Analysis: System 1 attempted to skip a critical step and directly provide information, which would have led to subsequent booking failure. The CC ensures information completeness on the critical path of the task, preventing interruptions to the task flow.
Case 4: CC Prevents a “Domain/Slot Confusion” Error
  • Background: A multi-domain scenario where the user wants to take a taxi from the hotel they are staying at.
  • Belief State:
  • System 1’s Incorrect Action: request(hotel, departure) with confidence $p_{t}^{S1}=0.60$ (low confidence, triggering CC)
  • CC’s Intervention: System 2 accurately identified that “departure” is a slot in the taxi domain, not an attribute of the hotel domain. System 2 corrected the action to: request(taxi, departure).
  • Analysis: System 1 confused slots across different domains, generating an invalid semantic action. Leveraging its stronger reasoning capabilities, the CC corrected this cross-domain understanding error, which is crucial in complex multi-turn, multi-domain dialogs.
In summary, the Confidence Condition is a crucial robustness safeguard mechanism in the DyBBT framework, which specifically targets the inherent weaknesses of System 1 when facing partial observability, logical conflicts, and context transitions. These errors are not only common but also fatal in nature. Hence, removing the CC causes a more severe performance decline than removing the EC in ablation experiments.
E.4 Supplementary Analysis Figures
@@
This subsection provides all supplementary figures supporting the main text analysis in Sec. 5, which offer intuitive data support for the discussions:
@@
  • Fig. 5: Heatmap of visitation frequency in the cognitive state space $\mathcal{C}$, illustrating the structured exploration strategy of the meta-controller across dialog phases.
@@
  • Fig. 5: Analysis of meta-controller decisions, showing the rate of System 2 invocation across dialog progress and the proportion of triggers from each condition.
@@
  • Fig. 7: Demonstrates the improvement of System 1 through knowledge distillation and the corresponding reduction in System 2 invocation over training.
@@
  • Fig. 7: Compares the empirical cumulative regret of DyBBT against the theoretical upper bound derived under simplifying assumptions.
@@
E.5 Hyperparameter Sensitivity Analysis
A key concern is the sensitivity of DyBBT’s performance to the meta-controller’s hyperparameters: the exploration threshold $\tau$, the confidence threshold $\kappa$, and the number of bins used to discretize the cognitive state space $\mathcal{C}$. We conducted a comprehensive grid search over $\tau\in\{0.5,1.0,1.5,2.0\}$, $\kappa\in\{0.5,0.6,0.7,0.8,0.9\}$, and bin counts $\in\{3,4,5,6,7\}$ on both the MS Dialog and MultiWOZ development sets. Performance is measured by the success rate (%), and the results are visualized in Fig. 8.